At 18 feet above ground level, Ryan steps onto an escalator that descends at a constant rate of 1.5 feet per second. Ryan's distance from the ground is given by the following formula where d is the distance from the ground, in feet,and n is the number of seconds he has been on the escalator.

d=18-1.5n
In how many seconds after Ryan steps onto the escalator will he be 8.25 feet from the ground?

Answer by rfer(16322)

9.75=18-1.5n
9.75-18=-1.5n
-8.25=-1.5n
n=5.5 sec